Dear colleagues,
Here is a summary of today's activites at CSBF.
- There was no power stoppage during the overnight muon run again.
- The CPT with CDM, which consists of 4 power modes, completed tonight.
- The remaining LPT of sub-detectors for the first CPT (GSE power supply) started from morning and ended at about 2 PM with an updated manual.
During the LPT, the data flow to the UMD SOC started but stopped around noon due to a rain outside at CSBF. No more connection until the end of the CPT. JiHye got more experience with this test especially on the data relay server setup at the cdps3. Plan to have an SOC expert training next week. Candidates are Dr. Kim, Opher, Ki-Chun and Brian. Other undergraduates can join the training also.
- Updated manuals will be uploaded on the blackboard before Monday.
- For the next 3 CPTs with different power modes (battery, battery being charged, and maximum power mode), CAL/tSCD/bSCD/CER pedestal runs in addition to the background muon runs in flight mode were carried out to test commanding from the Client PC and noise measurements in different power modes. They were all working OK w/o any problems.
- Manuals were updated based on today's experiences with changes in the cdaq software and parameters.
- The old PYL box shipped back to PSU for a repair. It will become a spare box.
- S3 threshold studied with different threshold values and the results are on the black board. With the correct parameters being changed, the S3Lo and ZLo were changed as expected.
- CD threshold issue related with the high ZHi rate problem solved. It was due to the wrong connector wiring and the bad connector at the MXR box. With Nick/Isaac's advice, the CD threshold wires moved into the connector pin #1 and #2 where TCD threshold control wires are in #3 and #4. After that, the CD threshold worked as expected. The measurements results are also on the same excel file as the above study. Also the CD signals and connector picture were uploaded on the black board as a reference.
- Prashant prepared cdps5 as a data relay machine in the local network. So the data taken in boram.creamnet will be copied to the cdps5 at the same time during the CSBF integration. This is very useful and would be better to be a part of future setups at CSBF and ICE.
- JiHye installed an event display program prepared by HoSeok at the cdps5. The same program is at UMD also.
- Debugging of the reduced hsk data could not be tried today by Prashant due to the CPT.
- SangEun continue working on the SIP interface coding before and after the CPT work.
Here are things to do.
- Reduced hsk debugging.
- Chercam HSK problem debugging.
- TCD 12 V HSK attenuation circuit installation.
- Mechanical integration starts from tomorrow by Fritz. The power of the instrument will be off during the mechanical work.
- Update of manuals including CPT.
- Copying and backup of newly taken integration data into cdps3 at UMD and client PCs at CSBF.
- Adding another Ethernet card to cdps5 for a dual connection to local network and to outside.
Please let us know if you have any questions/comments/suggestions.
Cheers, MooHyun.
MooHyun, July 24, 2008



